OpenStudy (anonymous):

HELPHELPHELP! Part 1: Solve the following equation using one of the methods listed below. Show all work. (4 points) Part 2: Using complete sentences, explain why you chose the method you used. (2 points) Equation: 0 = x2 + 6x + 3 Methods: Completing the Square Factoring Quadratic Formula Graphing

OpenStudy (anonymous):

completing the square is easier if you move the 3 to the other side of the equation. half of the 6x term and then square. Then add it to BOTH sides. Then the equation will factor to (x-3)^2=(9-3) then solve x-3=6 and you'll have both answers for x. Completing the square is a form of factoring, quadratic formula would yield the same thing but is considered "brute force" as it will give the answer regardless, typically in situations where the function won't factor. Graphing is easier as well, simply look for the points where y=0, but remember to not trust the calculator too much either. Calculus will show you why the calculator is actually wrong more times than it's right.